# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v1

@@ 6.1.2. GM fivefolds containing a fixed GM fourfold
According to [34, Section 5.2] and [21], there is a $\mathbb{P}^{1}$-fibration $\pi_{Y}\colon F_{qs}(Y)\twoheadrightarrow\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ and we have a commutative diagram
@@
where $F_{qs}(Y)$ is a connected component of the Hilbert scheme of quadric surfaces in $Y$ such that $S\cap X$ is a conic for each $[S]\in F_{qs}(Y)$ [^8] A quadric surface $S\subset Y$ is in $F_{qs}(Y)$ if and only if $S\subset\operatorname{Gr}(2,V_{4})$ for a $4$-dimensional subspace $V_{4}\subset V_{5}$. It is called a $\tau$-quadric in [15, Section 7.3].. Using this relation, the following result is obtained in [34, Proposition 5.6], which is an analog of Proposition 6.4.
@@
is of rank one. And from [18, Theorem 5.2(2)], we know that
@@
We can also demonstrate that $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is not injective as follows. If $i$ is an embedding, then by Theorem 6.7 and 6.13, $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\subset\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an atomic embedded Lagrangian. However, by [49, Equation (7.11)], we have
@@
which contradicts $\chi(\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2})=2\chi(\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2})=384$ by [23, Proposition 1.10].
@@ 6.2. Double EPW cubes
is constructed in [18, Theorem 5.7], such that $\widetilde{\mathsf{Z}}^{\geq 3}_{A(X)}$ is a smooth connected projective threefold when $X$ is general.
@@
Moreover, in [24, Section 8], the authors construct a rational map
@@
where $\pi_{Y}$ is induced by the projection functor, and $r$ is the morphism in Theorem 3.7. Moreover, $\pi_{Y}$ maps birationally onto a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ by [24, Theorem C.13], and $r$ is an embedding when restricted to this component by [24, Lemma 8.20].
In particular, denoting by $L_{Y}$ the image of $r\circ\pi_{Y}$, it follows that $L_{X}$ is isomorphic to a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ and $L_{Y}\subset M^{X}_{\sigma_{X}}(1,-1)\cong\widetilde{\mathsf{Z}}^{\geq 2}_{A(X)}$ is a Lagrangian subvariety. Moreover, $L_{Y}$ is smooth by [24, Theorem C.13, Corollary C.14] and $\mathrm{Hilb}^{3t+1}_{Y}\twoheadrightarrow L_{Y}$ is a divisorial contraction. Thus, Theorem 5.8 can be applied in this case, yielding:
@@ 7.1. The relation to the relative Jacobian fibration of GM fivefolds
First, we show that $s$ is injective, which is equivalent to showing that for any two different points $[Y],[Y^{\prime}]\in\mathcal{B}_{X}$, we have $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\neq\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$. Indeed, if $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}=\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$, then we have $\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}=\mathsf{Y}^{\geq 2}_{A(Y^{\prime})^{\perp}}\subset\mathsf{Y}_{A(X)^{\perp}}$. However, by [19, Proposition 2.5], we have a canonical identification $A(Y)=A(Y^{\prime})$, which implies that $Y=Y^{\prime}$ by [16, Propoition 3.14].
@@
Note that $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are reflexive sheaves by [29, Corollary 1.2]. Let $U\subset\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the smooth locus and $V\subset\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the pullback of $U$. Hence, $V\cong U$ and $\mathcal{N}_{1}|_{V}\cong\mathcal{N}_{2}|_{U}$. From the fact that $\dim(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\setminus V)=0$ by Lemma 6.2 and $\mathcal{N}_{2}$ is reflexive, we have
@@
Since $\mathcal{H}$ is smooth along $s(\mathcal{B}_{X})$, according to [22, Theorem 8.1], $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ is equipped with a holomorphic symplectic two-form and the relative Picard fibration
@@
is a Lagrangian fibration. In fact, while [22, Theorem 8.1] is stated only for embedded Lagrangians, it can also be applied in our case by Lemma 7.1. This is because the proof of [22, Theorem 8.1] involves a local study, as explained in [33, Section 3.1].
We denote by $\mathcal{M}$ the component of stable sheaves on $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ containing $i_{*}\mathcal{L}$, where $[Y]\in\mathcal{B}_{X}$ and $\mathcal{L}\in\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$. Note that $i_{*}\mathcal{L}$ is stable because it is a rank one torsion-free sheaf supported on the image of the normalization map $i$. Then $\mathcal{M}$ has an open dense subscheme $\mathcal{M}^{0}$ isomorphic to $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ such that the support morphism $\mathcal{M}^{0}\to\mathcal{B}_{X}$ is isomorphic to the relative Picard fibration via the identification of $\mathcal{B}_{X}$ as an open subscheme of $\mathcal{H}$ by Lemma 7.1. Thus, $\mathcal{M}^{0}$ also carries a holomorphic symplectic structure. This strategy is explained in [49, Section 7.4].
On the other hand, according to [19, Theorem 1.1], there is a canonical isomorphism of principally polarized abelian varieties
@@
for any $[Y]\in\mathcal{B}_{X}$. Thus, $\mathrm{J}(Y)\cong\mathrm{Alb}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})\cong\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$ and the fiber of $\mathcal{M}^{0}\to\mathcal{B}_{X}$ over $[Y]\in\mathcal{B}_{X}$ is canonically isomorphic to $\mathrm{J}(Y)$. This situation is very similar to the case of cubic fourfolds (cf. [22, Example 8.22] and [48, 33]), and we hope that the techniques in [48] can be applied to the Gushel–Mukai case to construct a suitable compactification $\overline{\mathcal{M}}$ of $\mathcal{M}^{0}\to\mathcal{B}_{X}$.
@@ 7.2.1. LLSvS eightfolds
Let $X$ be a cubic fourfold not containing a plane and $j\colon Y\hookrightarrow X$ be a smooth hyperplane section. Let $M_{3}(X)$ be the Hilbert scheme of generalized twisted cubics on $X$, which is a smooth projective tenfold by [45, Theorem 4.7].
@@
On the other hand, we have the moduli interpretation $\mathsf{Z}_{X}\cong M^{X}_{\sigma_{X}}(2,1)$ (see [47]) and $\mathsf{Z}_{Y}\cong M^{Y}_{\sigma_{Y}}(2,1)$ (see [3] or [24, Appendix B.2]). In particular, if we denote by $i\colon\mathsf{Z}_{Y}\hookrightarrow\mathsf{Z}_{X}$ the Lagrangian embedding constructed in [63], the moduli construction is shown to be compatible with $i$ by [24, Corollary B.7], as the following commutative diagram
@@
induced by the Stein factorization of the natural morphism $M_{3}(X)\rightarrow\operatorname{Gr}(4,V_{6})$, which maps a twisted cubic to its generating $\mathbb{P}^{3}$. Moreover, $\mathsf{Z}^{\prime}_{X}$ is smooth by [45, Theorem 4.7]. Similarly, there exists a smooth subvariety $\mathsf{Z}^{\prime}_{Y}\subset\mathsf{Z}^{\prime}_{X}$ induced by $M_{3}(Y)\subset M_{3}(X)$, and $f_{X}$ induces a morphism $f_{Y}\colon\mathsf{Z}^{\prime}_{Y}\to\operatorname{Gr}(4,V_{5})$, which maps a twisted cubic on $Y$ to its generating $\mathbb{P}^{3}$ as well. Therefore, it is straightforward to check that there is a Cartesian diagram
@@
According to [45, Theorem 4.11], there is an embedding $X\hookrightarrow\mathsf{Z}_{X}$ such that a point $[C]\in M_{3}(X)$ maps to $p\in X\subset\mathsf{Z}_{X}$ if and only if $C$ has $p$ as an embedded point. Moreover, the blow-up of $\mathsf{Z}_{X}$ along $X$ is $\mathsf{Z}_{X}^{\prime}$. We denote by $D\subset\mathsf{Z}_{X}^{\prime}$ the exceptional divisor, and let $\phi\colon\mathsf{Z}_{X}^{\prime}\to\mathsf{Z}_{X}$ be the blow-up morphism. As any two twisted cubics with embedded points in $Y$ have the same embedded point if and only if they span the same $\mathbb{P}^{3}$, we have an identification $\mathsf{Z}_{Y}^{\prime}=\mathsf{Z}_{Y}$. Then we obtain a commutative diagram
@@
By [45, pp.38], there exists an ample line bundle $L$ on $\mathsf{Z}_{X}$ such that we have an isomorphism $\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(D)\otimes f^{*}_{X}\operatorname{\mathcal{O}}_{\operatorname{Gr}(4,V_{6})}(1)\cong\phi^{*}L$. Moreover, according to [45, pp.37], we have $\omega_{\mathsf{Z}^{\prime}_{X}}\cong\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(3D)$, which implies
@@
then $D|_{\mathsf{Z}_{Y}}=0\in\mathrm{H}^{2}(\mathsf{Z}_{Y},\mathbb{Q})$. However, this is impossible since $\mathcal{O}_{Y}(D|_{Y})\cong\operatorname{\mathcal{O}}_{Y}(-1)$ by [45, Proposition 4.5]. Therefore, we see $c_{1}(\mathsf{Z}_{Y})\notin\operatorname{im}(i^{*})$ and $\mathsf{Z}_{Y}\subset\mathsf{Z}_{X}$ is not atomic by Theorem 4.10. This finishes the proof of (1).
@@ 7.2.2. Fano variety of lines
Besides this construction, the authors of [33] consider another Lagrangian surface of $F(X)$ arising from a cubic fivefold. Let $Z$ be a cubic fivefold, where $X$ is a hyperplane section of $Z$. We consider the Fano variety of planes of $Z$, denoted by $F_{2}(Z)$. When $X$ and $Z$ are general, $F_{2}(Z)$ is a smooth surface and there is a morphism $i\colon F_{2}(Z)\to F(X)$ such that $F_{2}(Z)$ is a connected immersed Lagrangian submanifold of $F(X)$ (cf. [33, Proposition 4]). Using a standard deformation argument of relative Hilbert schemes, one can show that $i^{*}\colon\mathrm{H}^{2}(F(X),\mathbb{Q})\to\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$ is of rank one, i.e. these Lagrangians deform in codimension one. However, we have the following result.
@@
We assume that $X=Z\cap\mathbb{P}(V_{6})\subset\mathbb{P}(V_{7})$. Let $\mathcal{T}^{\prime}$ be the pull-back of the rank two tautological subbundle on $\operatorname{Gr}(2,V_{6})$ via the natural embedding $F(X)\hookrightarrow\operatorname{Gr}(2,V_{6})$ and $\mathcal{W}:=V_{6}\otimes\operatorname{\mathcal{O}}_{F(X)}$. In [33, Lemma 6], it is shown that $i$ can be decomposed as
@@
If we denote by $\alpha:=c_{1}(\operatorname{\mathcal{O}}_{\mathbb{P}}(1))$ and $\beta:=c_{1}(\mathcal{T}^{\vee})$, then we get $c_{1}(\mathcal{K})=10\alpha+3\beta$ and $c_{1}(\mathbb{P})=4\alpha+\beta$. Hence, $c_{1}(F_{2}(Z))=14t^{*}\alpha+4t^{*}\beta$. However, as $h_{2}=t^{*}\beta$ is the pull-back of the class of the Plücker line bundle on $F(X)$, we have $0\neq t^{*}\beta\in\operatorname{im}(i^{*})$ and $\operatorname{im}(i^{*})$ is spanned by $t^{*}\beta$. Thus, as $c_{1}(F_{2}(Z))\in\operatorname{im}(i^{*})$, we have $t^{*}\alpha=0\in\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$. Since the fundamental class of $F_{2}(Z)$ in $\mathbb{P}$ is $63\alpha^{4}c_{2}(\mathcal{T}^{\vee})$, we obtain $063\alpha^{5}c_{2}(\mathcal{T}^{\vee})=\alpha.t_{*}[F_{2}(Z)]=t_{*}t^{*}\alpha=$ 0 . However, as computed in [33, Lemma 6], we have $\alpha^{5}=-\beta.\alpha^{4}$. This implies that
@@
Next, we prove (2). Since $\mathcal{S}/\mathcal{T}=\operatorname{\mathcal{O}}_{\mathbb{P}}(-1)$, we have $h_{1}-h_{2}=t^{*}\alpha$. Therefore, (2) follows from the description of $c_{1}(F_{2}(Z))$, Proposition 4.11, and Remark 4.12. ∎

### v2

@@ 6.1.2. GM fivefolds containing a fixed GM fourfold
According to [36, Section 5.2] and [22], there is a $\mathbb{P}^{1}$-fibration $\pi_{Y}\colon F_{qs}(Y)\twoheadrightarrow\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ and we have a commutative diagram
@@
where $F_{qs}(Y)$ is a connected component of the Hilbert scheme of quadric surfaces in $Y$ such that $S\cap X$ is a conic for each $[S]\in F_{qs}(Y)$ [^8] A quadric surface $S\subset Y$ is in $F_{qs}(Y)$ if and only if $S\subset\operatorname{Gr}(2,V_{4})$ for a $4$-dimensional subspace $V_{4}\subset V_{5}$. It is called a $\tau$-quadric in [16, Section 7.3].. Using this relation, the following result is obtained in [36, Proposition 5.6], which is an analog of Proposition 6.4.
@@
is of rank one. And from [19, Theorem 5.2(2)], we know that
@@
We can also demonstrate that $i\colon\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\to\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is not injective as follows. If $i$ is an embedding, then by Theorem 6.7 and 6.13, $\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}\subset\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ is an atomic embedded Lagrangian. However, by [54, Equation (7.11)], we have
@@
which contradicts $\chi(\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2})=2\chi(\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2})=384$ by [24, Proposition 1.10].
@@ 6.2. Double EPW cubes
is constructed in [19, Theorem 5.7], such that $\widetilde{\mathsf{Z}}^{\geq 3}_{A(X)}$ is a smooth connected projective threefold when $X$ is general.
@@
Moreover, in [25, Section 8], the authors construct a rational map
@@
where $\pi_{Y}$ is induced by the projection functor, and $r$ is the morphism in Theorem 3.7. Moreover, $\pi_{Y}$ maps birationally onto a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ by [25, Theorem C.13], and $r$ is an embedding when restricted to this component by [25, Lemma 8.20].
In particular, denoting by $L_{Y}$ the image of $r\circ\pi_{Y}$, it follows that $L_{X}$ is isomorphic to a connected component of $M_{\sigma_{Y}}^{Y}(1,-1)$ and $L_{Y}\subset M^{X}_{\sigma_{X}}(1,-1)\cong\widetilde{\mathsf{Z}}^{\geq 2}_{A(X)}$ is a Lagrangian subvariety. Moreover, $L_{Y}$ is smooth by [25, Theorem C.13, Corollary C.14] and $\mathrm{Hilb}^{3t+1}_{Y}\twoheadrightarrow L_{Y}$ is a divisorial contraction. Thus, Theorem 5.8 can be applied in this case, yielding:
@@ 7.1. The relation to the relative Jacobian fibration of GM fivefolds
First, we show that $s$ is injective, which is equivalent to showing that for any two different points $[Y],[Y^{\prime}]\in\mathcal{B}_{X}$, we have $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\neq\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$. Indeed, if $\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}=\overline{\mathsf{Y}}^{\geq 2}_{A(Y^{\prime})^{\perp}}$, then we have $\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}=\mathsf{Y}^{\geq 2}_{A(Y^{\prime})^{\perp}}\subset\mathsf{Y}_{A(X)^{\perp}}$. However, by [20, Proposition 2.5], we have a canonical identification $A(Y)=A(Y^{\prime})$, which implies that $Y=Y^{\prime}$ by [17, Propoition 3.14].
@@
Note that $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ are reflexive sheaves by [30, Corollary 1.2]. Let $U\subset\overline{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the smooth locus and $V\subset\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}$ be the pullback of $U$. Hence, $V\cong U$ and $\mathcal{N}_{1}|_{V}\cong\mathcal{N}_{2}|_{U}$. From the fact that $\dim(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}}\setminus V)=0$ by Lemma 6.2 and $\mathcal{N}_{2}$ is reflexive, we have
@@
Since $\mathcal{H}$ is smooth along $s(\mathcal{B}_{X})$, according to [23, Theorem 8.1], $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ is equipped with a holomorphic symplectic two-form and the relative Picard fibration
@@
is a Lagrangian fibration. In fact, while [23, Theorem 8.1] is stated only for embedded Lagrangians, it can also be applied in our case by Lemma 7.1. This is because the proof of [23, Theorem 8.1] involves a local study, as explained in [35, Section 3.1].
We denote by $\mathcal{M}$ the component of stable sheaves on $\widetilde{\mathsf{Y}}_{A(X)^{\perp}}$ containing $i_{*}\mathcal{L}$, where $[Y]\in\mathcal{B}_{X}$ and $\mathcal{L}\in\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$. Note that $i_{*}\mathcal{L}$ is stable because it is a rank one torsion-free sheaf supported on the image of the normalization map $i$. Then $\mathcal{M}$ has an open dense subscheme $\mathcal{M}^{0}$ isomorphic to $\operatorname{Pic}^{0}(\widetilde{\mathscr{Y}}^{\geq 2}/\mathcal{B}_{X})$ such that the support morphism $\mathcal{M}^{0}\to\mathcal{B}_{X}$ is isomorphic to the relative Picard fibration via the identification of $\mathcal{B}_{X}$ as an open subscheme of $\mathcal{H}$ by Lemma 7.1. Thus, $\mathcal{M}^{0}$ also carries a holomorphic symplectic structure. This strategy is explained in [54, Section 7.4].
On the other hand, according to [20, Theorem 1.1], there is a canonical isomorphism of principally polarized abelian varieties
@@
for any $[Y]\in\mathcal{B}_{X}$. Thus, $\mathrm{J}(Y)\cong\mathrm{Alb}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})\cong\operatorname{Pic}^{0}(\widetilde{\mathsf{Y}}^{\geq 2}_{A(Y)^{\perp}})$ and the fiber of $\mathcal{M}^{0}\to\mathcal{B}_{X}$ over $[Y]\in\mathcal{B}_{X}$ is canonically isomorphic to $\mathrm{J}(Y)$. This situation is very similar to the case of cubic fourfolds (cf. [23, Example 8.22] and [51, 35]), and we hope that the techniques in [51] can be applied to the Gushel–Mukai case to construct a suitable compactification $\overline{\mathcal{M}}$ of $\mathcal{M}^{0}\to\mathcal{B}_{X}$.
@@ 7.2.1. LLSvS eightfolds
Let $X$ be a cubic fourfold not containing a plane and $j\colon Y\hookrightarrow X$ be a smooth hyperplane section. Let $M_{3}(X)$ be the Hilbert scheme of generalized twisted cubics on $X$, which is a smooth projective tenfold by [48, Theorem 4.7].
@@
On the other hand, we have the moduli interpretation $\mathsf{Z}_{X}\cong M^{X}_{\sigma_{X}}(2,1)$ (see [50]) and $\mathsf{Z}_{Y}\cong M^{Y}_{\sigma_{Y}}(2,1)$ (see [3] or [25, Appendix B.2]). In particular, if we denote by $i\colon\mathsf{Z}_{Y}\hookrightarrow\mathsf{Z}_{X}$ the Lagrangian embedding constructed in [70], the moduli construction is shown to be compatible with $i$ by [25, Corollary B.7], as the following commutative diagram
@@
induced by the Stein factorization of the natural morphism $M_{3}(X)\rightarrow\operatorname{Gr}(4,V_{6})$, which maps a twisted cubic to its generating $\mathbb{P}^{3}$. Moreover, $\mathsf{Z}^{\prime}_{X}$ is smooth by [48, Theorem 4.7]. Similarly, there exists a smooth subvariety $\mathsf{Z}^{\prime}_{Y}\subset\mathsf{Z}^{\prime}_{X}$ induced by $M_{3}(Y)\subset M_{3}(X)$, and $f_{X}$ induces a morphism $f_{Y}\colon\mathsf{Z}^{\prime}_{Y}\to\operatorname{Gr}(4,V_{5})$, which maps a twisted cubic on $Y$ to its generating $\mathbb{P}^{3}$ as well. Therefore, it is straightforward to check that there is a Cartesian diagram
@@
According to [48, Theorem 4.11], there is an embedding $X\hookrightarrow\mathsf{Z}_{X}$ such that a point $[C]\in M_{3}(X)$ maps to $p\in X\subset\mathsf{Z}_{X}$ if and only if $C$ has $p$ as an embedded point. Moreover, the blow-up of $\mathsf{Z}_{X}$ along $X$ is $\mathsf{Z}_{X}^{\prime}$. We denote by $D\subset\mathsf{Z}_{X}^{\prime}$ the exceptional divisor, and let $\phi\colon\mathsf{Z}_{X}^{\prime}\to\mathsf{Z}_{X}$ be the blow-up morphism. As any two twisted cubics with embedded points in $Y$ have the same embedded point if and only if they span the same $\mathbb{P}^{3}$, we have an identification $\mathsf{Z}_{Y}^{\prime}=\mathsf{Z}_{Y}$. Then we obtain a commutative diagram
@@
By [48, pp.38], there exists an ample line bundle $L$ on $\mathsf{Z}_{X}$ such that we have an isomorphism $\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(D)\otimes f^{*}_{X}\operatorname{\mathcal{O}}_{\operatorname{Gr}(4,V_{6})}(1)\cong\phi^{*}L$. Moreover, according to [48, pp.37], we have $\omega_{\mathsf{Z}^{\prime}_{X}}\cong\operatorname{\mathcal{O}}_{\mathsf{Z}_{X}^{\prime}}(3D)$, which implies
@@
then $D|_{\mathsf{Z}_{Y}}=0\in\mathrm{H}^{2}(\mathsf{Z}_{Y},\mathbb{Q})$. However, this is impossible since $\mathcal{O}_{Y}(D|_{Y})\cong\operatorname{\mathcal{O}}_{Y}(-1)$ by [48, Proposition 4.5]. Therefore, we see $c_{1}(\mathsf{Z}_{Y})\notin\operatorname{im}(i^{*})$ and $\mathsf{Z}_{Y}\subset\mathsf{Z}_{X}$ is not atomic by Theorem 4.10. This finishes the proof of (1).
@@
Applying the argument in Theorem 1.2 to the atomic sheaf in Proposition 7.3(2), we have the following result.
Corollary 7.4.
[…]
@@ 7.2.2. Fano variety of lines
Besides this construction, the authors of [35] consider another Lagrangian surface of $F(X)$ arising from a cubic fivefold. Let $Z$ be a cubic fivefold, where $X$ is a hyperplane section of $Z$. We consider the Fano variety of planes of $Z$, denoted by $F_{2}(Z)$. When $X$ and $Z$ are general, $F_{2}(Z)$ is a smooth surface and there is a morphism $i\colon F_{2}(Z)\to F(X)$ such that $F_{2}(Z)$ is a connected immersed Lagrangian submanifold of $F(X)$ (cf. [35, Proposition 4]). Using a standard deformation argument of relative Hilbert schemes, one can show that $i^{*}\colon\mathrm{H}^{2}(F(X),\mathbb{Q})\to\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$ is of rank one, i.e. these Lagrangians deform in codimension one. However, we have the following result.
@@
We assume that $X=Z\cap\mathbb{P}(V_{6})\subset\mathbb{P}(V_{7})$. Let $\mathcal{T}^{\prime}$ be the pull-back of the rank two tautological subbundle on $\operatorname{Gr}(2,V_{6})$ via the natural embedding $F(X)\hookrightarrow\operatorname{Gr}(2,V_{6})$ and $\mathcal{W}:=V_{6}\otimes\operatorname{\mathcal{O}}_{F(X)}$. In [35, Lemma 6], it is shown that $i$ can be decomposed as
@@
If we denote by $\alpha:=c_{1}(\operatorname{\mathcal{O}}_{\mathbb{P}}(1))$ and $\beta:=c_{1}(\mathcal{T}^{\vee})$, then we get $c_{1}(\mathcal{K})=10\alpha+3\beta$ and $c_{1}(\mathbb{P})=4\alpha+\beta$. Hence, $c_{1}(F_{2}(Z))=14t^{*}\alpha+4t^{*}\beta$. However, as $h_{2}=t^{*}\beta$ is the pull-back of the class of the Plücker line bundle on $F(X)$, we have $0\neq t^{*}\beta\in\operatorname{im}(i^{*})$ and $\operatorname{im}(i^{*})$ is spanned by $t^{*}\beta$. Thus, as $c_{1}(F_{2}(Z))\in\operatorname{im}(i^{*})$, we have $t^{*}\alpha=0\in\mathrm{H}^{2}(F_{2}(Z),\mathbb{Q})$. Since the fundamental class of $F_{2}(Z)$ in $\mathbb{P}$ is $63\alpha^{4}c_{2}(\mathcal{T}^{\vee})$, we obtain $063\alpha^{5}c_{2}(\mathcal{T}^{\vee})=\alpha.t_{*}[F_{2}(Z)]=t_{*}t^{*}\alpha=$ 0 . However, as computed in [35, Lemma 6], we have $\alpha^{5}=-\beta.\alpha^{4}$. This implies that
@@
Next, we prove (2). Since $\mathcal{S}/\mathcal{T}=\operatorname{\mathcal{O}}_{\mathbb{P}}(-1)$, we have $h_{1}-h_{2}=t^{*}\alpha$. Therefore, (2) follows from the description of $c_{1}(F_{2}(Z))$, Proposition 4.11, and Remark 4.12.
Applying the argument in Theorem 1.2 to the atomic sheaf in Proposition 7.5(2), we have the following result.
Corollary 7.6.
[…]